OpenStudy (anonymous):

andy has $14 more than jack together they have $56 how much dose each boy have

OpenStudy (anonymous):

@bernardo2131341 Let's call Andy 'a' and Jack 'j'. We know that a = j +14 and also that a + j = 56 so we can sub the first equation into the second. (j +14) + j = 56 2j + 14 = 56 2j = 42 j = 21 Now that we have a value for j, we can sub this back into the first equation. a = 21 + 14 a = 35 So Jack has 21 and Andy has 35.